### v2

# Improved bounds for randomized Schatten norm estimation of numerically low-rank matrices

Abstract
In this work, we analyze the variance of a stochastic estimator for computing Schatten norms of matrices. The estimator extracts information from a single sketch of the matrix, that is, the product of the matrix with a few standard Gaussian random vectors. While this estimator has been proposed and used in the literature before, the existing variance bounds are often pessimistic. Our work provides a sharper upper bound on the variance and we also give estimates of the variance that work well for numerically low-rank matrices. Our theoretical findings are supported by numerical experiments, demonstrating that the new bounds are significantly tighter than the existing ones.
Keywords.
Schatten norm, randomized estimator, variance bound, trace estimator, numerically low-rank matrices, spectrum moments estimation
MSC code.
65F35, 62J10, 68W20
1 Introduction
Given a matrix $A\in\mathbb{R}^{m\times n}$, we consider the problem of numerically approximating its Schatten-$2p$ norm
$\|A\|_{2p}=\left(\sigma_{1}^{2p}+\sigma_{2}^{2p}+\cdots+\sigma_{\min\{n,m\}}^{2p}\right)^{\frac{1}{2p}},$
where $\sigma_{1},\sigma_{2},\ldots,\sigma_{\min\{n,m\}}$ are the singular values of $A$ and $p$ is a positive integer. While the Schatten-$2p$ norm can be readily computed from the singular values of the matrix, computing the singular values themselves generally requires $\mathcal{O}(nm\min\{n,m\})$ time and is infeasible for large-scale matrices.
Several randomized methods have been developed to compute an approximation of the Schatten norm with cheaper operations, trading some accuracy for speed. A class of such methods is based on the Hutchinson trace estimator [5], a randomized algorithm that approximates the trace of a square matrix $B\in\mathbb{R}^{n\times n}$ by averaging a few quadratic forms $\omega_{i}^{T}B\omega_{i}$, where $\omega_{i}$’s are isotropic random vectors of length $n$. The Hutchinson estimator is often used when $B$ is a matrix function [1, 12] and relates to the Schatten-$2p$ norm estimation problem via the formula
$\|A\|_{2p}^{2p}=\mathrm{trace}\left((A^{T}A)^{p}\right).$ (1)
A variety of variance reduction techniques has been developed, for instance the combination of Hutchinson trace estimator with randomized low-rank approximation; see, e.g., [3, 9, 11]. Algorithms specifically tailored to Schatten-$p$ norm estimation have been discussed in [2, 4]. When computing quadratic forms involving $(A^{T}A)^{p}$, the matrix $A$ needs to be accessed at least $p$ times. This number can be slightly reduced if, instead of computing exact quadratic forms, these are approximated with Chebyshev polynomials as in [2, 4].
If, instead, we are only allowed to access the matrix $A$ once, we can use the Schatten-$2p$ norm estimator proposed by Kong and Valiant [6], an unbiased estimator for $\|A\|_{2p}^{2p}$. The problem considered in [6], restricted to the Gaussian setting, is to extract information on the eigenvalues $\lambda_{1},\ldots,\lambda_{n}$ of a covariance matrix $B=A^{T}A$ given access to some samples drawn from the multivariate Gaussian distribution corresponding to $B$. This is approached via the methods of moments, which requires estimating the $p$-th moment of the spectrum of $B$ defined as $\sum_{i=1}^{n}\lambda_{i}^{p}$. In the language of randomized numerical linear algebra, Kong and Valiant’s algorithm [6, Algorithm 1] estimates Schatten-$2p$ norms of a matrix $A$ in the so-called streaming setting. The samples drawn from the distributions correspond to a sketch of $A$, that is, to a matrix $Y:=A\Omega$ where $\Omega\in\mathbb{R}^{n\times k}$ consists of i.i.d. standard Gaussian entries, for some $k\ll n$. Throughout the paper, we denote the estimator for $\|A\|_{2p}^{2p}$ from [6] as $\hat{\theta}_{2p}$ and summarize its construction procedure in Algorithm 1. In Algorithm 1, line 2 uses MATLAB function notation triu($\cdot$,1) to indicate the strictly upper triangular part of a matrix.
Algorithm 1 Estimator for Schatten-$2p$ norm from [6]
0:  Sketch $Y=A\Omega\in\mathbb{C}^{m\times k}$ and integer $p$
0:  Unbiased estimator $\hat{\theta}_{2p}$ for $\|A\|_{2p}^{2p}$
1:  Compute $Z=Y^{T}Y$.
2:  Define $T=\mathrm{triu}(Z,1)$ the strictly upper triangular part of the matrix $Z$.
3:  Compute $\hat{\theta}_{2p}=\binom{k}{p}^{-1}\mathrm{trace}(T^{p-1}Z)$
It was proved in [6, Proposition 4] that the variance of $\hat{\theta}_{2p}$ with a Gaussian sketching matrix $\Omega$ satisfies the following bound:
$\mathrm{Var}(\hat{\theta}_{2p})\leq 2^{12p}p^{6p}3^{p}\max\left\{\frac{n^{p-2}}{k^{p}},\frac{1}{k},\frac{n^{\frac{1}{2}-\frac{1}{p}}}{k}\right\}\|A\|_{2p}^{4p}.$ (2)
However, this bound usually severely overestimates the true variance of $\hat{\theta}_{2p}$. As an illustrative example, Figure 1 plots the variance for $\hat{\theta}_{6}$ (estimator of the $6$-th power of the Schatten-$6$ norm) on a $10\times 10$ matrix with singular values $1,1/2,1/3,\ldots,1/10$, together with the bound (2). The bound overestimates the variance by several orders of magnitude.
Figure 1: Comparison between true variance of $\hat{\theta}_{6}$ and the variance bound (2) for a $10\times 10$ matrix $A$.
We will show in Theorem 3 that, with a Gaussian sketching matrix $\Omega$, the variance of $\hat{\theta}_{2p}$ has first-order expansion
$\mathrm{Var}(\hat{\theta}_{2p})=\frac{2p^{2}}{k}\|A\|_{4p}^{4p}+\mathcal{O}\left(\frac{1}{k^{2}}\right).$ (3)
Remark 1.
We remark that the variance of $\hat{\theta}_{2p}$ is usually worse than the variance of estimators based on the Hutchinson trace estimator, though both are unbiased. Specifically, since the variance of the vanilla Hutchinson trace estimator [5] for a matrix $B$ with Gaussian random vectors is $2\|B\|_{F}^{2}$, applying it to (1) gives an unbiased estimator of $\|A\|_{2p}^{2p}$ with variance
$\frac{2p}{k}\|A^{2p}\|_{F}^{2}=\frac{2p}{k}\|A\|_{4p}^{4p},$
which is smaller than (3). We remark that, in order to use the Hutchinson trace estimator, the matrix $A$ needs to be accessed more than once (to apply powers of the matrix to a vector). When the only available information about the matrix $A$ is its sketch $Y$, to the best of the authors’ knowledge, the estimator $\hat{\theta}_{2p}$ proposed in [6] is the only method to estimate the Schatten-$2p$ norms of $A$.
Contributions.
In this paper, we focus on the streaming setting and we are interested in better understanding the variance of $\hat{\theta}_{2p}$ and its behavior on matrices with strongly decaying singular values (i.e., numerically low-rank matrices). We derive an improved bound on the variance of $\hat{\theta}_{2p}$ in the case of a Gaussian sketch. Our bounds leverage the rotational invariance of the Gaussian distribution to simplify the analysis of the estimator, and the results we obtain only depend on the singular values of $A$, highlighting a better behavior in the case of rapid singular value decay. Numerical results show that this bound is relatively tight for matrices with rapid singular value decay using moderately large sketch size $k$. Moreover, in all cases our variance bound is more precise than the previous bound (2). We provide a first-order and second-order expansion of the variance of $\hat{\theta}_{2p}$ in $1/k$. The experiments verify that our first-order and second-order estimates capture the behavior of $\hat{\theta}_{2p}$ for matrices that exhibit a strong singular value decay, for moderate values of $k$.
Outline.
Section 2 reviews the building blocks of $\hat{\theta}_{2p}$. Section 3 presents our main results on the variance of $\hat{\theta}_{2p}$. In Section 4, we illustrate the performance of our bounds and estimates on a variety of examples with/without singular value decay. Our conclusions are summarized in Section 5.
2 Notations and review of estimator $\hat{\theta}_{2p}$
This short section defines the notations throughout the paper and introduces a mathematically equivalent expression of $\hat{\theta}_{2p}$, which will be useful in our analysis later. Given a positive integer $k$, we denote $[k]=\{1,2,\ldots,k\}$. Given another positive integer $p\leq k$, a $p$-cycle is a sequence of $p$ distinct integers $\tau=(i_{1},i_{2},\ldots,i_{p})\in[k]^{p}$. When a matrix $Z\in\mathbb{R}^{k\times k}$ is further given, every $p$-cycle $\tau$ defines a product
$Z_{\tau}=\prod_{\ell=1}^{k}Z_{i_{\ell},i_{\ell+1}},$
with the convention that $i_{k+1}\equiv i_{1}$ for simpler notation. With this notation, we are ready to state one key idea towards the construction of $\hat{\theta}_{2p}$.
Proposition 2 ([7] and [6, Fact 2]).
For any $p$-cycle $\tau=(i_{1},\ldots,i_{p})$ with $i_{\ell}\in[k]$, a real matrix $A\in\mathbb{R}^{m\times n}$, and an $n\times k$ random matrix $\Omega=(\omega_{j,\ell})$ with i.i.d. entries with mean 00 and variance $1$,
$\mathbb{E}[(\Omega^{T}A^{T}A\Omega)_{\tau}]=\mathrm{trace}((A^{T}A)^{p})=\|A\|_{2p}^{2p},$
where the expectation is over the randomness of $\Omega$.
Setting $Y=A\Omega$, Proposition 2 says that $(Y^{T}Y)_{\tau}$ is an unbiased estimate of $\|A\|_{2p}^{2p}$ for any (fixed) $p$-cycle $\tau$. A natural approach to reduce the variance of this estimator is to average over all possible $p$-cycles. However, this strategy poses significant computational challenges, as naively iterating over all $\frac{k!}{(k-p)!}$ possible cycles becomes computationally infeasible, even for moderately large values of $k$. A surprising linear algebra fact is that the average over all increasing $p$-cycles, that is, $p$-cycles such that $i_{1}<i_{1}<\ldots<i_{p}$, can be efficiently computed by simple linear algebra operations, and this defines $\hat{\theta}_{2p}$ presented in Algorithm 1:
$\hat{\theta}_{2p}=\binom{k}{p}^{-1}\sum_{1\leq i_{1}<\cdots<i_{p}\leq k}(\Omega^{T}A^{T}A\Omega)_{\tau}=\binom{k}{p}^{-1}\mathrm{trace}(T^{p-1}Z),$ (4)
where $Z$ and $T$ are defined as in Algorithm 1. The proof for last identity in (4) can be found in [6, Lemma 1]. We note that Proposition 2 together with the relation (4) guarantee the unbiasedness of $\hat{\theta}_{2p}$.
3 Improved variance estimates for $\hat{\theta}_{2p}$
This section presents our main results regarding the variance of $\hat{\theta}_{2p}$ and the analysis. All the theorems are stated in Section 3.1 while all the proofs are deferred to Section 3.2.
3.1 Main results
Our first result provides the first order expansion in $1/k$ for the variance of $\hat{\theta}_{2p}$ with $k$ being the sketch size (Theorem 3).
Theorem 3.
Let $\hat{\theta}_{2p}$ be the unbiased estimator for $\|A\|_{2p}^{2p}$ given in Algorithm 1. Then the variance of $\hat{\theta}_{2p}$ is
$\mathrm{Var}(\hat{\theta}_{2p})=\frac{2p^{2}\|A\|_{4p}^{4p}}{k}+\mathcal{O}\left(\frac{1}{k^{2}}\right).$
Moreover, the constant in $\mathcal{O}(\cdot)$ is a symmetric function of the singular values of $A$.
The next theorem (Theorem 4) provides an upper bound for the variance of a scaling of $\hat{\theta}_{2p}$. The bound is a function of $k$, $p$, and Schatten norms of $A$, highlighting lower variance for numerically low-rank matrices.
Theorem 4.
Let $\hat{\Theta}_{2p}\coloneqq\binom{k}{p}\hat{\theta}_{2p}$. Then the variance of $\hat{\Theta}_{2p}$ is bounded by
$\mathrm{Var}(\hat{\Theta}_{2p})$ $\leq-\binom{k}{p}^{2}\|A\|_{2p}^{4p}+\binom{k}{2p}\binom{2p}{p}\|A\|_{2p}^{4p}+\binom{k}{2p-1}\binom{2p-1}{p}p\left(\|A\|_{2p}^{4p}+2\|A\|_{4p}^{4p}\right)$
$\quad+\binom{k}{2p-2}\binom{2p-2}{p}\binom{p}{2}\left(6\|A\|_{4p}^{4p}+3\|A\|_{4}^{4}\|A\|_{4p-4}^{4p-4}\right)$
$\quad+\sum_{r=3}^{p}\binom{k}{2p-r}\binom{2p-r}{p}\binom{p}{r}\sum_{\ell=1}^{r}c(r,\ell)\|A\|_{4}^{4(\ell-1)}\|A\|_{4p-4(\ell-1)}^{4p-4(\ell-1)},$
where
$c(r,\ell)=\begin{cases}3^{r}&\text{if }\ell=1;\\ 3^{r-2}(2^{r+1}-1)&\text{if }\ell=2;\\ 3^{r-\ell}\ell^{r}&\text{otherwise.}\end{cases}$
Dividing the bound for $\mathrm{Var}(\hat{\Theta}_{2p})$ in Theorem 4 by $\binom{k}{p}^{2}$ immediately provides an upper bound for $\mathrm{Var}(\hat{\theta}_{2p})$. The second-order estimate for the variance of $\hat{\theta}_{2p}$ presented in Corollary 5 below follows from Theorem 4.
Corollary 5.
The variance of $\hat{\theta}_{2p}$ satisfies
$\mathrm{Var}(\hat{\theta}_{2p})\leq\frac{2p^{2}\|A\|_{4p}^{4p}}{k}+\frac{p^{2}(p-1)^{2}}{k^{2}}\left(\|A\|_{4p}^{4p}+\frac{3}{2}\|A\|_{4}^{4}\|A\|_{4p-4}^{4p-4}-\frac{1}{2}\|A\|_{2p}^{4p}\right)+\mathcal{O}\left(\frac{1}{k^{3}}\right).$
3.2 Proofs of the main results
The proofs of our main results are separated into four subsections. Section 3.2.1 breaks down the analysis for $\mathrm{Var}(\hat{\theta}_{2p})$ into several subtasks. Section 3.2.2 presents the proof of first-order result (Theorem 3). Section 3.2.3 and Section 3.2.4 are devoted for the upper bound (Theorem 4) and second-order estimate (Corollary 5).
3.2.1 Breaking Var($\hat{\theta}_{2p}$) into mathematically manageable pieces
Let $A$ be an $m\times n$ deterministic matrix and $\Omega\in\mathbb{R}^{n\times k}$ a random matrix with i.i.d. $N(0,1)$ entries. Considering the full singular value decomposition $A=U\Sigma V^{T}$, where $U\in\mathbb{R}^{m\times m}$, $\Sigma\in\mathbb{R}^{m\times n}$, and $V\in\mathbb{R}^{n\times n}$, the entries of $\widetilde{\Omega}:=V^{T}\Omega$ are also i.i.d. $N(0,1)$ random variables, due to the rotation invariance of the standard Gaussian distribution. Therefore, the matrix $Z$ in line 1 of Algorithm 1 can be written as
$Z=(A\Omega)^{T}(A\Omega)=\widetilde{\Omega}^{T}\Sigma^{T}U^{T}U\Sigma\widetilde{\Omega}=\widetilde{\Omega}^{T}\Sigma^{T}\Sigma\widetilde{\Omega}.$
Denoting
$\Sigma^{T}\Sigma=\begin{bmatrix}\sigma_{1}^{2}&&&&\\ &\sigma_{2}^{2}&&&\\ &&\ddots&&\\ &&&\ddots&\\ &&&&\sigma_{n}^{2}\end{bmatrix}\quad\text{ and }\quad\widetilde{\Omega}=\begin{bmatrix}\omega_{1,1}&\omega_{1,2}&\cdots&\omega_{1,k}\\ \omega_{2,1}&\omega_{2,2}&\cdots&\omega_{2,k}\\ \vdots&\vdots&&\vdots\\ \vdots&\vdots&&\vdots\\ \omega_{n,1}&\omega_{n,2}&\cdots&\omega_{n,k}\end{bmatrix}$
with the convention that $\sigma_{\min\{n,m\}+1}=\ldots=\sigma_{\max\{n,m\}}=0$ and using the relation (4), the estimator $\hat{\theta}_{2p}$ from Algorithm 1 takes the form:
$\hat{\theta}_{2p}$ $=\binom{k}{p}^{-1}\sum_{i_{1}<\cdots<i_{p}}(\sigma_{1}^{2}\omega_{1,i_{1}}\omega_{1,i_{2}}+\cdots+\sigma_{n}^{2}\omega_{n,i_{1}}\omega_{n,i_{2}})\cdots(\sigma_{1}^{2}\omega_{1,i_{p}}\omega_{1,i_{1}}+\cdots+\sigma_{n}^{2}\omega_{n,i_{p}}\omega_{n,i_{1}})$
$=\binom{k}{p}^{-1}\sum_{i_{1}<\cdots<i_{p}}\prod_{h=1}^{p}\left(\sum_{t=1}^{n}\sigma_{t}^{2}\omega_{t,i_{h}}\omega_{t,i_{h+1}}\right),$
with the indices $i_{1},\ldots,i_{p}$ ranging from $1$ to $k$ and with the convention that $i_{p+1}\equiv i_{1}$. For ease of notation, define $\hat{\Theta}_{2p}:=\binom{k}{p}\hat{\theta}_{2p}$ so that
$\mathrm{Var}(\hat{\theta}_{2p})=\binom{k}{p}^{-2}\left(\mathbb{E}[\hat{\Theta}_{2p}^{2}]-\mathbb{E}[\hat{\Theta}_{2p}]^{2}\right)=\binom{k}{p}^{-2}\mathbb{E}[\hat{\Theta}_{2p}^{2}]-\|A\|_{2p}^{4p},$ (5)
where the second equality follows because $\hat{\theta}_{2p}$ is an unbiased estimator of $\|A\|_{2p}^{2p}$. It remains to estimate/bound the second moment of $\hat{\Theta}_{2p}$. Note that
$\hat{\Theta}_{2p}^{2}=\sum_{i_{1}<\cdots<i_{p}\atop j_{1}<\cdots<j_{p}}\prod_{h=1}^{p}\left[\left(\sum_{t=1}^{n}\sigma_{t}^{2}\omega_{t,i_{h}}\omega_{t,i_{h+1}}\right)\left(\sum_{t=1}^{n}\sigma_{t}^{2}\omega_{t,j_{h}}\omega_{t,j_{h+1}}\right)\right],$ (6)
where the indices $i_{1},\ldots,i_{p}$ and $j_{1},\ldots,j_{p}$ range from $1$ to $k$. Estimating or bounding the variance of $\hat{\theta}_{2p}$ involves estimating or bounding the expectation of each term in the outer sum from (6). For ease of notation, we denote
$f_{i_{1},\ldots,j_{p}}:=\prod_{h=1}^{p}\left[\left(\sum_{t=1}^{n}\sigma_{t}^{2}\omega_{t,i_{h}}\omega_{t,i_{h+1}}\right)\left(\sum_{t=1}^{n}\sigma_{t}^{2}\omega_{t,j_{h}}\omega_{t,j_{h+1}}\right)\right]$ (7)
and let $F_{r}$ be the sum of $\mathbb{E}[f_{i_{1},\ldots,j_{p}}]$ over the choices of $i_{1}<\cdots<i_{p}$ and $j_{1}<\cdots<j_{p}$ with $r$ repeated indices:
$F_{r}:=\sum_{2p\text{-tuples }i_{1}<\cdots<i_{p},j_{1}<\cdots<j_{p}\text{ among }\atop\text{ which we have }r\text{ repeated indices}}\mathbb{E}[f_{i_{1},\ldots,j_{p}}].$ (8)
Then, we can split the second moment of $\hat{\Theta}_{2p}$ as
$\mathbb{E}[\hat{\Theta}_{2p}^{2}]=\sum_{r=0}^{p}\left(\sum_{2p\text{-tuples }i_{1}<\cdots<i_{p},j_{1}<\cdots<j_{p}\text{ among }\atop\text{ which we have }r\text{ repeated indices}}\mathbb{E}[f_{i_{1},\ldots,j_{p}}]\right)=\sum_{r=0}^{p}F_{r}.$ (9)
Our analysis follows by estimating or bounding each $F_{r}$. Lemmas 6 and 7 below describes the contribution of each $F_{r}$ to $\mathbb{E}[\hat{\Theta}_{2p}^{2}]$ in terms of the order of $k$.
Lemma 6.
The number of $2p$-tuples $i_{1},\ldots,i_{p},j_{1},\ldots,j_{p}$ such that $i_{1}<\cdots<i_{p}$, $j_{1}<\cdots<j_{p}$, and among $i_{1},\ldots,j_{p}$ there are $r$ repeated indices $(0\leq r\leq p)$ is $\binom{k}{2p-r}\binom{2p-r}{p}\binom{p}{r}$.
Proof.
The set $\{i_{1},\ldots,j_{p}\}$ contains $2p-r$ distinct values from the set $\{1,\ldots,k\}$, so there are $\binom{k}{2p-r}$ possibilities for this. There are now $\binom{2p-r}{p}$ ways of assigning $i_{1},\ldots,i_{p}$ to some of the chosen indices. Finally, there are $\binom{p}{r}$ ways of choosing which of these $p$ indices are repeated among $j_{1},\ldots,j_{p}$. Once we made the decision, the assignment is uniquely determined by the fact that $i_{1}<\cdots<i_{p}$ and $j_{1}<\cdots<j_{p}$. ∎
Lemma 7.
For each $0\leq r\leq p$, the quantity $F_{r}$ defined in (8), as a polynomial in $k$, has degree at most $\mathcal{O}(k^{2p-r})$. Hence, the second moment $\mathbb{E}[\hat{\Theta}_{2p}^{2}]$, as a polynomial in $k$, has degree at most $2p$.
Proof.
Since the distribution of $\omega_{i,j}$ is independent of $k$ for any index $(i,j)$, so is the expectation $\mathbb{E}[f_{i_{1},\ldots,j_{p}}]$. By Lemma 6, $F_{r}$ grows (as a function in $k$) at the same rate as $\binom{k}{2p-r}\binom{2p-r}{p}\binom{p}{r}$, which is a polynomial in $k$ of degree $2p-r$. ∎
3.2.2 First-order analysis
Recall from (5) and (9), the variance of $\hat{\theta}_{2p}$ is
$\mathrm{Var}(\hat{\theta}_{2p})=\binom{k}{p}^{-2}\mathbb{E}[\hat{\Theta}_{2p}^{2}]-\|A\|_{2p}^{4p}=\binom{k}{p}^{-2}\sum_{r=0}^{p}F_{r}-\|A\|_{2p}^{4p}.$ (10)
Since $\binom{k}{p}^{2}$ is a polynomial (in $k$) of degree $2p$ and, by Lemma 7, $F_{r}$ is a polynomial in $k$ of degree $2p-r$, it suffices to consider $F_{0}$ and $F_{1}$ in the first-order analysis. Therefore, we are interested in the sum of $\mathbb{E}[f_{i_{1},\ldots,j_{p}}]$ over the choices of $i_{1}<\cdots<i_{p}$ and $j_{1}<\cdots<j_{p}$ with either zero or one repeated index. Expanding the product of the two summations in (7), the expectation $\mathbb{E}[f_{i_{1},\ldots,j_{p}}]$ can be further decomposed into the sum of $n^{2p}$ expectations:
$\mathbb{E}[f_{i_{1},\ldots,j_{p}}]=\sum_{(t_{1},\ldots,s_{p})\in[n]^{2p}}\mathbb{E}\left[Y_{i_{1},\ldots,j_{p}}^{t_{1},\ldots,s_{p}}\right],$ (11)
where the random variable $Y_{i_{1},\ldots,j_{p}}^{t_{1},\ldots,s_{p}}$ takes the form
$Y_{i_{1},\ldots,j_{p}}^{t_{1},\ldots,s_{p}}:=\prod_{h=1}^{p}\sigma_{t_{h}}^{2}\omega_{t_{h},i_{h}}\omega_{t_{h},i_{h+1}}\sigma_{s_{h}}^{2}\omega_{s_{h},j_{h}}\omega_{s_{h},j_{h+1}}.$ (12)
Observe that, if any of the $\omega$-variables appearing in (12) has an odd exponent, the expectation of $Y_{i_{1},\ldots,j_{p}}^{t_{1},\ldots,s_{p}}$ becomes zero since all odd moments of standard Gaussian random variables are zero. This observation gives the following proposition.
Proposition 8.
Suppose $i_{h}$ (resp. $j_{h}$), $h\in\{1,2,\ldots,p\}$, is a non-repeated index among $i_{1}<\cdots<i_{p}$ and $j_{1}<\cdots<j_{p}$. Then a necessary condition for $\mathbb{E}[Y_{i_{1},\ldots,j_{p}}^{t_{1},\ldots,s_{p}}]$ to be nonzero is that $t_{h-1}=t_{h}$ (resp. $s_{h-1}=s_{h}$), with the convention that $t_{0}\equiv t_{p}$ and $s_{0}\equiv s_{p}$ for simpler notation.
Proof.
The product of the random variables in (12) is
$\omega_{t_{1},i_{1}}\omega_{t_{1},i_{2}}\omega_{t_{2},i_{2}}\omega_{t_{2},i_{3}}\cdots\omega_{t_{p},i_{p}}\omega_{t_{p},i_{1}}\omega_{s_{1},j_{1}}\omega_{s_{1},j_{2}}\omega_{s_{2},j_{2}}\omega_{s_{2},j_{3}}\cdots\omega_{s_{p},j_{p}}\omega_{s_{p},j_{1}}.$ (13)
If $Y_{i_{1},\ldots,j_{p}}^{t_{1},\ldots,s_{p}}$ has nonzero expectation, each $\omega$-variable in (13) must have even exponent. In particular, there must be a second copy of $\omega_{t_{h},i_{h}}$. This can happen only if $\omega_{t_{h},i_{h}}=\omega_{t_{h-1},i_{h}}$ since $i_{h}$ is different from any other index in $\{i_{1},\ldots,j_{p}\}\setminus\{i_{h}\}$ (i.e., $i_{h}$ is non-repeated index). Hence, $t_{h-1}=t_{h}$. The argument for the $s$-indices is the same. ∎
Leveraging Proposition 8, we can compute $\mathbb{E}[f_{i_{1},\ldots,j_{p}}]$ in the cases where there are either no repeated indices or only one repeated index among $i_{1},\ldots,j_{p}$.
Lemma 9.
If there is no repeated index among $i_{1}<\cdots<i_{p}$ and $j_{1}<\cdots<j_{p}$, then
$\mathbb{E}[f_{i_{1},\ldots,j_{p}}]=\|A\|_{2p}^{4p}.$
Proof.
Consider a single nonzero term $\mathbb{E}[Y_{i_{1},\ldots,j_{p}}^{t_{1},\ldots,s_{p}}]$ on the right-hand-side of (11). Then $t_{1}=\ldots=t_{p}$ and $s_{1}=\ldots=s_{p}$ by Proposition 8 since there is no repeated index among $i_{1}<\cdots<i_{p}$ and $j_{1}<\cdots<j_{p}$. In this case, we have
$\mathbb{E}[Y_{i_{1},\ldots,j_{p}}^{t_{1},\ldots,s_{p}}]=\mathbb{E}\left[\sigma_{t_{1}}^{2p}\sigma_{s_{1}}^{2p}\prod_{h=1}^{p}\omega_{t_{1},i_{h}}^{2}\omega_{s_{1},j_{h}}^{2}\right]=\sigma_{t_{1}}^{2p}\sigma_{s_{1}}^{2p},$
where the last equality follows from the fact that the $\omega_{t_{1},i_{h}}$ and $\omega_{s_{1},j_{h}}$ are independent random variables since there is no repeated index among $i_{1},\ldots,j_{p}$. To obtain $\mathbb{E}[f_{i_{1},\ldots,j_{p}}]$ we need to sum over all possible choices of $t_{1}$ and $s_{1}$, each of which can be anything in $\{1,\ldots,n\}$. Therefore,
$\mathbb{E}[f_{i_{1},\ldots,j_{p}}]=\sum_{t=1}^{n}\sum_{s=1}^{n}\sigma_{t}^{2p}\sigma_{s}^{2p}=\|A\|_{2p}^{4p}.\qed$
Lemma 10.
If there is exactly one repeated index among $i_{1}<\cdots<i_{p}$ and $j_{1}<\cdots<j_{p}$, then
$\mathbb{E}[f_{i_{1},\ldots,j_{p}}]=\|A\|_{2p}^{4p}+2\|A\|_{4p}^{4p}.$
Proof.
Let $\mathbb{E}[Y_{i_{1},\ldots,j_{p}}^{t_{1},\ldots,s_{p}}]$ be a nonzero term on the right-hand-side of (11). Without loss of generality, let us assume that the repeated index is $i_{1}$. Then Proposition 8 guarantees that $t_{h}=t_{h-1}$ for $h=2,\ldots,p$, so that $t_{1}=\ldots=t_{p}$. The same argument also proves $s_{1}=\ldots=s_{p}$.
Now, given $t_{1}=\ldots=t_{p}$ and $s_{1}=\ldots=s_{p}$, we consider two disjoint subcases:
  • If $t_{1}\neq s_{1}$, then
    $\mathbb{E}[Y_{i_{1},\ldots,j_{p}}^{t_{1},\ldots,s_{p}}]=\mathbb{E}\left[\sigma_{t_{1}}^{2p}\sigma_{s_{1}}^{2p}\prod_{h=1}^{p}\omega_{t_{1},i_{h}}^{2}\omega_{s_{1},j_{h}}^{2}\right]=\sigma_{t_{1}}^{2p}\sigma_{s_{1}}^{2p},$
    where the last equality follows from the fact that the $\omega_{t_{1},i_{h}}$ and $\omega_{s_{1},j_{h}}$ are independent random variables since $t_{1}\neq s_{1}$.
  • If $t_{1}=s_{1}$, then the fact that $i_{1},\ldots,i_{p}$ and $j_{1},\ldots,j_{p}$ have exactly one repeated index means that the product (12) has exactly one $N(0,1)$ random variable to the fourth power and all the others are squares of $N(0,1)$ random variables. Therefore,
    $\mathbb{E}[Y_{i_{1},\ldots,j_{p}}^{t_{1},\ldots,s_{p}}]=\mathbb{E}\left[\sigma_{t_{1}}^{2p}\sigma_{t_{1}}^{2p}\prod_{h=1}^{p}\omega_{t_{1},i_{h}}^{2}\omega_{t_{1},j_{h}}^{2}\right]=3\sigma_{t_{1}}^{4p},$
    where the factor of $3$ comes from the single fourth moment of $N(0,1)$.
To obtain $\mathbb{E}[f_{i_{1},\ldots,j_{p}}]$, we sum over all possible choices of $t_{1}$ and $s_{1}$, each of which can be anything in $\{1,\ldots,n\}$. Therefore,
$\mathbb{E}[f_{i_{1},\ldots,j_{p}}]=3\sum_{t=1}^{n}\sigma_{t}^{4p}+\sum_{t\neq s}\sigma_{t}^{2p}\sigma_{s}^{2p}=3\|A\|_{4p}^{4p}+\left(\|A\|_{2p}^{4p}-\|A\|_{4p}^{4p}\right)=\|A\|_{2p}^{4p}+2\|A\|_{4p}^{4p}.\qed$
Given Lemmas 9 and 10, we are ready to compute the first-order expansion of $\mathrm{Var}(\hat{\theta}_{2p})$ in $1/k$ presented in Theorem 3.
Proof of Theorem 3.
Combining Lemmas 6, 9 and 10, we obtain
$F_{0}$ $=\binom{k}{2p}\binom{2p}{p}\|A\|_{2p}^{4p}=\frac{k^{2p}-p(2p-1)k^{2p-1}}{(p!)^{2}}\|A\|_{2p}^{4p}+\mathcal{O}(k^{2p-2}),$
$F_{1}$ $=\binom{k}{2p-1}\binom{2p-1}{p}p(\|A\|_{2p}^{4p}+2\|A\|_{4p}^{4p})$
$=\frac{k^{2p-1}}{[(p-1)!]^{2}}(\|A\|_{2p}^{4p}+2\|A\|_{4p}^{4p})+\mathcal{O}(k^{2p-2}).$
All the constants in $\mathcal{O}(\cdot)$ involve only $p$, $\|A\|_{2p}$, and $\|A\|_{4p}$, so they are symmetric functions in singular values of $A$. Moreover, Lemma 7 guarantees
$\sum_{r=2}^{p}F_{r}=\mathcal{O}(k^{2p-2})$
and the constants in $\mathcal{O}(\cdot)$ are again symmetric functions in singular values of $A$ since each $\mathbb{E}[f_{i_{1},\ldots,j_{p}}]$ is. Combining all the above, the variance of $\hat{\theta}_{2p}$ becomes
$\mathrm{Var}(\hat{\theta}_{2p})$ $=\binom{k}{p}^{-2}\mathbb{E}[\hat{\Theta}_{2p}^{2}]-\|A\|_{2p}^{4p}$
$=\binom{k}{p}^{-2}\left[F_{0}+F_{1}+\sum_{r=2}^{p}F_{r}-\binom{k}{p}^{2}\|A\|_{2p}^{4p}\right]$
$=\frac{(p!)^{2}}{k^{2}(k-1)^{2}\cdots(k-p+1)^{2}}\left[\frac{k^{2p}-p(2p-1)k^{2p-1}}{(p!)^{2}}\|A\|_{2p}^{4p}\right.$
$\left.\qquad+\frac{k^{2p-1}}{[(p-1)!]^{2}}(\|A\|_{2p}^{4p}+2\|A\|_{4p}^{4p})-\frac{k^{2p}-p(p-1)k^{2p-1}}{(p!)^{2}}\|A\|_{2p}^{4p}+\mathcal{O}(k^{2p-2})\right]$
$=\frac{2p^{2}\|A\|_{2p}^{4p}}{k}+\mathcal{O}\left(\frac{1}{k^{2}}\right),$
where all the constants inside $\mathcal{O}(\cdot)$ are symmetric functions of singular values of $A$. ∎
3.2.3 Terms with at least 2 repeated indices
To obtain an upper bound for $\mathrm{Var}(\hat{\theta}_{2p})$, we need to further consider $i_{1}<\cdots<i_{p}$ and $j_{1}<\cdots<j_{p}$ with $r$ repeated indices and $r\geq 2$. Motivated by the decomposition in (11), we essentially want to count or bound the number of tuples $(t_{1},\ldots,s_{p})\in[n]^{2p}$ such that $Y_{i_{1},\ldots,j_{p}}^{t_{1},\ldots,s_{p}}$ has nonzero expectation. To better visualize the product
$Y_{i_{1},\ldots,j_{p}}^{t_{1},\ldots,s_{p}}=\prod_{h=1}^{p}\sigma_{t_{h}}^{2}\omega_{t_{h},i_{h}}\omega_{t_{h},i_{h+1}}\sigma_{s_{h}}^{2}\omega_{s_{h},j_{h}}\omega_{s_{h},j_{h+1}}$ (14)
for some fixed $i_{1},\ldots,i_{p},j_{1},\ldots,j_{p}$, let us consider two circles with $p$ points each, corresponding to $i_{1},\ldots,i_{p}$ and $j_{1},\ldots,j_{p}$, respectively. This is illustrated in Figure 2 for $p=8$. The $r$ repeated indices are denoted with an extra circle; in Figure 2 we have 3 repeated indices, that is, $i_{1}=j_{2}$, $i_{2}=j_{4}$, and $i_{5}=j_{8}$. Note that such pairings between repeated indices are unique since $i_{1}<\cdots<i_{p}$ and $j_{1}<\cdots<j_{p}$. Fixing the indices $i_{1},\ldots,j_{p}$, a single choice of $(t_{1},\ldots,s_{p})\in[n]^{2p}$ for $Y_{i_{1},\ldots,j_{p}}^{t_{1},\ldots,s_{p}}$ corresponds to assigning singular value $\sigma_{t_{h}}$ to the segment $(i_{h},i_{h+1})$, and assigning singular value $\sigma_{s_{h}}$ to the segment $(j_{h},j_{h+1})$.
$i_{1}$$i_{2}$$i_{3}$$i_{4}$$i_{5}$$i_{6}$$i_{7}$$i_{8}$
$j_{1}$$j_{2}$$j_{3}$$j_{4}$$j_{5}$$j_{6}$$j_{7}$$j_{8}$
Figure 2: Circle notation.
Let $i_{h}$ be a non-repeated index (corresponding to non-circled point in Figure 2). Proposition 8 asserts that the two segments $(i_{h-1},i_{h})$ and $(i_{h},i_{h+1})$ adjacent to point $i_{h}$ must be assigned to the same singular value; otherwise the expectation of $Y_{i_{1},\ldots,j_{p}}^{t_{1},\ldots,s_{p}}$ will vanish. To illustrate this fact, in Figure 2, we used the same color to highlight the segments that need to be assigned to the same singular value as a necessary condition for the expectation of the term (12) to be nonzero. In the following, we refer to the group of contiguous segments with the same color as an “arc”.
Consider a choice of indices $i_{1}<\cdots<i_{p}$ and $j_{1}<\cdots<j_{p}$ with $r$ repeated indices. Then, each circle has $r$ colored arcs. Moreover, there is a natural bijection from the colored arcs in $i$-circle to the color arcs in $j$-circle, where each colored arcs is mapped to the colored arc in the other circle with the same indices as endpoints. In our example, the pairings are $(i_{1},i_{2})$$(j_{2},j_{4})$, $(i_{2},i_{5})$$(j_{4},j_{8})$, and $(i_{5},i_{1})$$(j_{8},j_{2})$, where each arc is intended in a counter-clockwise direction. We denote the arcs by $a_{1}^{(i)},\ldots,a_{r}^{(i)},a_{1}^{(j)},\ldots,a_{r}^{(j)}$ and we will say that $a_{\ell}^{(i)}$ and $a_{\ell}^{(j)}$ are corresponding pairs of arcs for $\ell=1,\ldots,r$.
The next proposition limits the possible singular values assignment to $2r$ arcs in the two circles and is crucial for the proofs of Lemma 13 and Lemma 14.
Proposition 11.
Fix a choice of $i_{1}<\cdots<i_{p}$ and $j_{1}<\cdots<j_{p}$ with $r\geq 2$ repeated indices and a $2p$-tuple $(t_{1},\ldots,s_{p})\in[n]^{2p}$ such that $\mathbb{E}[Y_{i_{1},\ldots,j_{p}}^{t_{1},\ldots,s_{p}}]$ is nonzero. Consider the visualization for $Y_{i_{1},\ldots,j_{p}}^{t_{1},\ldots,s_{p}}$ as in Figure 2. Suppose the arcs $a_{1}^{(i)}$ and $a_{2}^{(i)}$ in the $i$-circle are assigned to singular values $\sigma_{g}$ and $\sigma_{h}$, respectively, with $g\neq h$. Then there are only two possible singular values assignments to their corresponding arcs $a_{1}^{(j)}$ and $a_{2}^{(j)}$ in the $j$-circle:
  1. (1)
    $a_{1}^{(j)}$ is assigned to $\sigma_{g}$ and $a_{2}^{(j)}$ is assigned to $\sigma_{h}$. Moreover, if this is the case, then for each $\ell=1,\ldots,r$, the singular values assigned to corresponding pair of arcs $a_{\ell}^{(i)}$ and $a_{\ell}^{(j)}$ must be the same. (Singular values assigned to different pairs could be different.)
  2. (2)
    $a_{1}^{(j)}$ is assigned to $\sigma_{h}$ and $a_{2}^{(j)}$ is assigned to $\sigma_{g}$. Moreover, if this is the case, then every arc $a_{\ell}^{(i)}$ in $i$-circle must be assigned to either $\sigma_{g}$ or $\sigma_{h}$; while the corresponding arc $a_{\ell}^{j}$ in $j$-circle must be assigned to the other singular value. That is, no singular values other than $\sigma_{g}$ and $\sigma_{h}$ can be present in this case.
Proof.
Let $i^{\ast}_{2}$ denote the common repeated index (common endpoint) for arcs $a_{1}^{(i)}$ and $a_{2}^{(i)}$. Note that $i^{\ast}_{2}$ is also the common endpoint of $a_{1}^{(j)}$ and $a_{2}^{(j)}$. Since $a_{1}^{(i)}$ is assigned to $\sigma_{g}$ and $a_{2}^{(i)}$ is assigned to $\sigma_{h}$, the product of the $\omega$-variables in (14) contains $\omega_{g,i_{2}^{\ast}}$ and $\omega_{h,i_{2}^{\ast}}$, which are different random variables since $g\neq h$. To have nonzero $\mathbb{E}[Y_{i_{1},\ldots,j_{p}}^{t_{1},\ldots,s_{p}}]$, there must be a second copy of both $\omega_{g,i_{2}^{\ast}}$ and $\omega_{h,i_{2}^{\ast}}$, and they must come from the segments joining the common endpoint of $a_{1}^{(j)}$ and $a_{2}^{(j)}$ in $j$-circle since the repeated index $i_{2}^{\ast}$ does not appear in the endpoint of any other segment in two circles. This means that one of the arc $a_{1}^{(j)}$ and $a_{2}^{(j)}$ has to be assigned to $\sigma_{g}$ and the other has to be assigned to $\sigma_{h}$, which proves the fact that there are only two possible singular values assignment to $a_{1}^{(j)}$ and $a_{2}^{(j)}$ in $j$-circle.
Now, consider case 1 and case 2 separately. Denote $i_{3}^{\ast}$ as the repeated index commonly appear at the endpoints of $a_{2}^{(i)}$, $a_{2}^{(j)}$, $a_{3}^{(i)}$, and $a_{3}^{(j)}$.
  1. (1)
    In this case, both $a_{2}^{(i)}$ and $a_{2}^{(j)}$ are assigned to $\sigma_{h}$. Such assignment contributes $\omega_{h,i_{3}^{\ast}}^{2}$ in the product of (14). To guarantee that the other two $\omega_{\bullet,i_{3}^{\ast}}$ from endpoint $i_{3}^{\ast}$ of $a_{3}^{(i)}$ and $a_{3}^{(j)}$ coincide (so that $\mathbb{E}[Y_{i_{1},\ldots,j_{p}}^{t_{1},\ldots,s_{p}}]$ is nonzero), the corresponding pair of arcs $a_{3}^{(i)}$ and $a_{3}^{(j)}$ must be assigned to the same singular value. Inductively, one can prove that, for every $\ell=1,\ldots,r$, the corresponding arcs $a_{\ell}^{(i)}$ and $a_{\ell}^{(j)}$ must be assigned to the same singular value.
  2. (2)
    In this case, $a_{2}^{(i)}$ is assigned to $\sigma_{h}$ and $a_{2}^{(j)}$ is assigned to $\sigma_{g}$. This assignment contributes $\omega_{h,i_{3}^{\ast}}\omega_{g,i_{3}^{\ast}}$ in the product of (14). To guarantee both $\omega_{h,i_{3}^{\ast}}$ and $\omega_{h,i_{3}^{\ast}}$ have even exponent in the product of (14), one of the arc $a_{3}^{(i)}$ and $a_{3}^{(j)}$ has to be assigned to $\sigma_{g}$ and the other has to be assigned to $\sigma_{h}$. The proof is complete by induction. ∎
Let us introduce one more notation which will be used later in our proofs. We assume each segment in the circles of Figure 2 has unit length. Denote $d_{\ell}^{(i)}$ the length of arc $a_{\ell}^{(i)}$ and $d_{\ell}^{(j)}$ the length of arc $a_{\ell}^{(j)}$, for $\ell=1,\ldots,r$. All lengths are at least $1$ and we have
$\sum_{\ell=1}^{r}d_{\ell}^{(i)}=\sum_{\ell=1}^{r}d_{\ell}^{(j)}=p.$
Our analysis will also use the following inequality between products of Schatten norms.
Lemma 12.
Let $2\leq c\leq d$ be two integers. Then $\|A\|_{c}^{c}\|A\|_{d}^{d}\leq\|A\|_{c-1}^{c-1}\|A\|_{d+1}^{d+1}$.
Proof.
We need to show that, for any sequence of nonnegative integers $\sigma_{1},\ldots,\sigma_{n}$,
$\sum\sigma_{i}^{c}\sum\sigma_{j}^{d}\leq\sum\sigma_{i}^{c-1}\sum\sigma_{j}^{d+1}.$
Note that there is a bijection between the terms $\sigma_{i}^{c}\sigma_{j}^{d}$ on the left-hand-side and the terms $\sigma_{i}^{c-1}\sigma_{j}^{d+1}$ on the right-hand-side. Therefore, it is sufficient to show that, for all $i,j\in\{1,\ldots,n\}$, we have
$\sigma_{i}^{c}\sigma_{j}^{d}+\sigma_{i}^{d}\sigma_{j}^{c}\leq\sigma_{i}^{c-1}\sigma_{j}^{d+1}+\sigma_{i}^{d+1}\sigma_{j}^{c-1}.$ (15)
We have that
$\sigma_{i}^{c-1}\sigma_{j}^{d+1}-\sigma_{i}^{c}\sigma_{j}^{d}-\sigma_{i}^{d}\sigma_{j}^{c}+\sigma_{i}^{d+1}\sigma_{j}^{c-1}=\sigma_{i}^{c-1}\sigma_{j}^{c-1}(\sigma_{j}-\sigma_{i})\left(\sigma_{j}^{d-c+1}-\sigma_{i}^{d-c+1}\right).$ (16)
Since $d\geq c$, we also have that $d-c+1\geq 0$, therefore $\sigma_{j}^{d-c+1}-\sigma_{i}^{d-c+1}$ has the same sign as $\sigma_{j}-\sigma_{i}$. Hence, (16) is always nonnegative, which implies (15) by rearranging the terms, and therefore implies the result of the lemma. ∎
Now, we are ready to bound $\mathbb{E}[f_{i_{1},\ldots,j_{p}}]$ when there are two or more repeated indices among $i_{1}<\cdots<i_{p}$ and $j_{1}<\cdots<j_{p}$.
Lemma 13.
If there are exactly two repeated indices among $i_{1}<\cdots<i_{p}$ and $j_{1}<\cdots<j_{p}$, then
$\mathbb{E}[f_{i_{1},\ldots,j_{p}}]\leq 6\|A\|_{4p}^{4p}+3\|A\|_{4}^{4}\|A\|_{4p-4}^{4p-4}.$
Proof.
Since there are two repeated indices, there are two arcs for each circle. Using our previous notation, they are $a_{1}^{(i)}$, $a_{2}^{(i)}$, $a_{1}^{(j)}$, and $a_{2}^{(j)}$. By Proposition 11, the nonzero expectation $\mathbb{E}[Y_{i_{1},\ldots,j_{p}}^{t_{1},\ldots,s_{p}}]$ on the right-hand-side of (11) corresponds to one of the following situations.
  • All arcs have the same singular value:
    $\mathbb{E}[Y_{i_{1},\ldots,j_{p}}^{t_{1},\ldots,s_{p}}]=\mathbb{E}\left[\sigma_{t_{1}}^{2p}\sigma_{t_{1}}^{2p}\prod_{h=1}^{p}\omega_{t_{1},i_{h}}^{2}\omega_{t_{1},j_{h}}^{2}\right]=9\sigma_{t_{1}}^{4p},$
    where the factor $9$ comes from the fact that two indices are repeated and therefore there are two fourth powers of i.i.d. $N(0,1)$ random variables.
  • The arcs in the $i$-circle have one singular value $\sigma_{t_{1}}$ and the arcs in the $j$-circle have another singular value $\sigma_{s_{1}}$:
    $\mathbb{E}[Y_{i_{1},\ldots,j_{p}}^{t_{1},\ldots,s_{p}}]=\mathbb{E}\left[\sigma_{t_{1}}^{2p}\sigma_{s_{1}}^{2p}\prod_{h=1}^{p}\omega_{t_{1},i_{h}}^{2}\omega_{s_{1},j_{h}}^{2}\right]=\sigma_{t_{1}}^{2p}\sigma_{s_{1}}^{2p}.$
  • Arcs $a_{1}^{(i)}$ and $a_{1}^{(j)}$ have the same singular value $\sigma_{t_{1}}$, and arcs $a_{2}^{(i)}$ and $a_{2}^{(j)}$ have the same singular value $\sigma_{t_{2}}$, with $t_{1}\neq t_{2}$:
    $\mathbb{E}[Y_{i_{1},\ldots,j_{p}}^{t_{1},\ldots,s_{p}}]=\mathbb{E}\left[\sigma_{t_{1}}^{2(d_{1}^{(i)}+d_{1}^{(j)})}\sigma_{t_{2}}^{2(d_{2}^{(i)}+d_{2}^{(j)})}\prod_{h=1}^{p}\omega_{t_{1},i_{h}}^{2}\omega_{s_{1},j_{h}}^{2}\right]=\sigma_{t_{1}}^{2(d_{1}^{(i)}+d_{1}^{(j)})}\sigma_{t_{2}}^{2(d_{2}^{(i)}+d_{2}^{(j)})}.$
  • Arcs $a_{1}^{(i)}$ and $a_{2}^{(j)}$ have the same singular value $\sigma_{t_{1}}$, and arcs $a_{2}^{(i)}$ and $a_{1}^{(j)}$ have the same singular value $\sigma_{t_{2}}$, with $t_{1}\neq t_{2}$:
    $\mathbb{E}[Y_{i_{1},\ldots,j_{p}}^{t_{1},\ldots,s_{p}}]=\mathbb{E}\left[\sigma_{t_{1}}^{2(d_{1}^{(i)}+d_{2}^{(j)})}\sigma_{t_{2}}^{2(d_{2}^{(i)}+d_{1}^{(j)})}\prod_{h=1}^{p}\omega_{t_{1},i_{h}}^{2}\omega_{s_{1},j_{h}}^{2}\right]=\sigma_{t_{1}}^{2(d_{1}^{(i)}+d_{2}^{(j)})}\sigma_{t_{2}}^{2(d_{2}^{(i)}+d_{1}^{(j)})}.$
Now we need to sum up all these contributions:
$\mathbb{E}[f_{i_{1},\ldots,j_{p}}]$ $=9\sum_{t=1}^{n}\sigma_{t}^{4p}+\sum_{t\neq s}\left(\sigma_{t}^{2p}\sigma_{s}^{2p}+\sigma_{t}^{2(d_{1}^{(i)}+d_{1}^{(j)})}\sigma_{s}^{2(d_{2}^{(i)}+d_{2}^{(j)})}+\sigma_{t}^{2(d_{1}^{(i)}+d_{2}^{(j)})}\sigma_{s}^{2(d_{2}^{(i)}+d_{1}^{(j)})}\right)$
$=9\|A\|_{4p}^{4p}+\left(\|A\|_{2p}^{4p}+\|A\|_{2(d_{1}^{(i)}+d_{1}^{(j)})}^{2(d_{1}^{(i)}+d_{1}^{(j)})}+\|A\|_{2(d_{1}^{(i)}+d_{2}^{(j)})}^{2(d_{1}^{(i)}+d_{2}^{(j)})}-3\|A\|_{4p}^{4p}\right)$
$\leq 6\|A\|_{4p}^{4p}+3\|A\|_{4}^{4}\|A\|_{4p-4}^{4p-4},$
where we used Lemma 12 (repeatedly) for the inequality at the end, since $2(d_{1}^{(i)}+d_{1}^{(j)})\geq 4$ and $2(d_{1}^{(i)}+d_{2}^{(j)})\geq 4$. ∎
Lemma 14.
If there are $r>2$ repeated indices among $i_{1}<\cdots<i_{p}$ and $j_{1}<\cdots<j_{p}$, then
$\mathbb{E}[f_{i_{1},\ldots,j_{p}}]\leq\sum_{\ell=1}^{r}c(r,\ell)\|A\|_{4}^{4(\ell-1)}\|A\|_{4p-4(\ell-1)}^{4p-4(\ell-1)}.$
where
$c(r,\ell)=\begin{cases}3^{r}&\text{if }\ell=1;\\ 3^{r-2}(2^{r+1}-1)&\text{if }\ell=2;\\ 3^{r-\ell}\ell^{r}&\text{otherwise.}\end{cases}$ (17)
Proof.
It suffices to consider the terms $Y_{i_{1},\ldots,j_{p}}^{t_{1},\ldots,s_{p}}$ with nonzero expectation. By Proposition 11, the maximum number of different colors in the two circles of Figure 2 (i.e., the maximum number of different singular values in the product (14)) is $r$. We separately consider the terms $Y_{i_{1},\ldots,j_{p}}^{t_{1},\ldots,s_{p}}$ in which there are $\ell$ distinct colors (singular values), for $1\leq\ell\leq r$.
  • Case $\ell=1$. All the arcs in the two circles have the same color $t$. Then,
    $\mathbb{E}[Y_{i_{1},\ldots,j_{p}}^{t_{1},\ldots,s_{p}}]=\mathbb{E}[Y_{i_{1},\ldots,j_{p}}^{t,\ldots,t}]=3^{r}\sigma_{t}^{4p},$
    where the coefficient $3^{r}$ comes from the fact that there are $r$ repeated indices and therefore $r$ i.i.d. $N(0,1)$ random variables that are to the fourth power in $Y_{i_{1},\ldots,j_{p}}^{t,\ldots,t}$.
  • Case $\ell=2$. Without loss of generality, assume the two colors are $t_{1}$ and $t_{2}$. First of all, the maximum number of fourth powers of random variables in the product (14) is $r-2$, which gives a coefficient of at most $3^{r-2}$ in front of the singular values. Due to Proposition 11, there are only two possible cases 1 and 2:
    1. (1)
      For each $\ell=1,\ldots,r$, the two corresponding arcs $a_{\ell}^{(i)}$ and $a_{\ell}^{(j)}$ have the same color. Moreover, not all $r$ pairs have the same color. There are $12^{r}-$ 1 color assignments of colors $t_{1}$ and $t_{2}$ that satisfy this constraint; each of them has a certain color pattern, denoted by “patt”, which uniquely determines the total length of arcs colored by $t_{1}$:
      $d(\text{patt})=\sum_{\ell\text{ such that }\atop a_{\ell}^{(i)}\text{has color}t_{1}}d_{\ell}^{(i)}+\sum_{\ell\text{ such that }\atop a_{\ell}^{(j)}\text{has color }t_{1}}d_{\ell}^{(j)}.$ (18)
      Note that $22\leq d(\text{patt})\leq 2p-$ 2 since not all pairs of arcs have same color. Then, the product of singular values in (14) is
      $\sigma_{t_{1}}^{2d(\text{patt})}\sigma_{t_{2}}^{4p-2d(\text{patt})}.$ (19)
    2. (2)
      For each $\ell=1,\ldots,r$, the two corresponding arcs $a_{\ell}^{(i)}$ and $a_{\ell}^{(j)}$ have different colors. There are $2^{r}$ color assignments of colors $t_{1}$ and $t_{2}$ that satisfy this constraint. Each color pattern again uniquely determines a number $d(\text{patt})$ in (18), and hence the the product of singular values in (14) is, again, (19).
    Then, we need to sum over all possible choices of $t_{1}$ and $t_{2}$, each of which can be anything in $\{1,\ldots,n\}$, so that
[…]
  • Case $3\leq\ell\leq r$. Denote the distinct colors by $t_{1},\ldots,t_{\ell}$. First of all, by Proposition 11, only case 1 can happen: this means that, for each $\ell=1,\ldots,r$, the two corresponding arcs $a_{\ell}^{(i)}$ and $a_{\ell}^{(j)}$ have the same color.
    We claim that the maximum number of fourth power of standard Gaussian random variables in product (14) is $r-\ell$. To see this, notice that we get a fourth power of the random variable $\omega_{\bullet,i^{\ast}}$ where $i^{\ast}$ is a repeated index (circled point) if and only if the two contiguous arcs joined by $i^{\ast}$ have the same color in both circles. For $r$ arcs in the circle with $\ell$ distinct colors, there are at most $r-\ell+1$ adjacent arcs that can have the same color. Hence, there are at most $r-\ell$ fourth power of standard Gaussian. This contributes to a coefficient at most $3^{r-\ell}$ in front of singular values in $\mathbb{E}[Y_{i_{1},\ldots,j_{p}}^{t_{1},\ldots,s_{p}}]$.
    Since we are under Proposition 11, case 1, the color of $a_{\ell}^{(i)}$ uniquely determines the color of $a_{\ell}^{(j)}$. It therefore suffices to determine the color pattern for $i$-circle. A valid color pattern needs to have all $\ell$ different colors in $i$-cycle. The total number of valid color patterns is bounded by $\ell^{r}$ (which is the number of ways to assign $\ell$ colors to $r$ arcs in $i$-circle).
    Now, given a valid color pattern “patt” of colors $t_{1},\ldots,t_{\ell}$ on $2r$ arcs in the two circles, it uniquely determines the total length of arcs colored by each $t_{q}$, $1\leq q\leq\ell$,
    $d_{q}(\text{patt})=\sum_{\ell\text{ such that }\atop a_{\ell}^{(i)}\text{has color }t_{q}}d_{\ell}^{(i)}+\sum_{\ell\text{ such that }\atop a_{\ell}^{(j)}\text{has color }t_{q}}d_{\ell}^{(j)}.$
    They satisfy the relations $d_{q}(\text{patt})\geq 2$ and $d_{1}(\text{patt})+\cdots+d_{\ell}(\text{patt})=2p$. With this color pattern, the expectation of product (14) is bounded above by $3^{r-\ell}\sigma_{t_{1}}^{2d_{1}(\text{patt})}\sigma_{t_{2}}^{2d_{2}(\text{patt})}\cdots\sigma_{t_{\ell}}^{2d_{\ell}(\text{patt})}$.
    Finally, summing over all possible color patterns and all possible distinct colors, we have
[…]
    $\leq 3^{r-\ell}\sum_{\text{valid color}\atop\text{pattern patt}}\sum_{\text{ distinct}\atop t_{1},\ldots,t_{\ell}}\sigma_{t_{1}}^{2d_{1}(\text{patt})}\sigma_{t_{2}}^{2d_{2}(\text{patt})}\cdots\sigma_{t_{\ell}}^{2d_{\ell}(\text{patt})}$
    $\leq 3^{r-\ell}\sum_{\text{valid color}\atop\text{pattern patt}}\prod_{q=1}^{\ell}\left(\sum_{t}\sigma_{t}^{2d_{q}(\text{patt})}\right)=3^{r-\ell}\sum_{\text{valid color}\atop\text{pattern patt}}\prod_{q=1}^{\ell}\|A\|_{2d_{q}(\text{patt})}^{2d_{q}(\text{patt})}$
    $\leq 3^{r-\ell}\sum_{\text{valid color}\atop\text{pattern patt}}\|A\|_{4}^{4(\ell-1)}\|A\|_{4p-4(\ell-1)}^{4p-4(\ell-1)}\leq 3^{r-\ell}\ell^{r}\|A\|_{4}^{4(\ell-1)}\|A\|_{4p-4(\ell-1)}^{4p-4(\ell-1)}.$
    where the second to the last inequality is obtained by repeatedly applying Lemma 12.
Summing all the contributions of $\mathbb{E}[Y_{i_{1},\ldots,j_{p}}^{t_{1},\ldots,s_{p}}]$ from the above cases, we have
$\mathbb{E}[f_{i_{1},\ldots,j_{p}}]\leq\sum_{\ell=1}^{r}c(r,\ell)\|A\|_{4}^{4(\ell-1)}\|A\|_{4p-4(\ell-1)}^{4p-4(\ell-1)}$
for the function $c(r,\ell)$ defined in the statement of the lemma. ∎
3.2.4 Second-order analysis
Now, we are ready to prove the bounds in Theorem 4 and Corollary 5.
Proof of Theorem 4.
Recalling (9) and the fact that $\hat{\Theta}_{2p}$ is an unbiased estimator for $\binom{k}{p}\|A\|_{2p}^{2p}$ and (9), we have
$\mathrm{Var}(\hat{\Theta}_{2p})=\mathbb{E}[\hat{\Theta}_{2p}^{2}]-\left(\mathbb{E}[\hat{\Theta}_{2p}]\right)^{2}=F_{0}+F_{1}+F_{2}+\sum_{r=3}^{p}F_{r}-\binom{k}{p}^{2}\|A\|_{2p}^{4p}.$
The desired bound in the statement of theorem immediately follows by applying Lemmas 6, 9, 10, 13 and 14 to obtain
$\mathbb{E}[\hat{\Theta}_{2p}^{2}]$ $=F_{0}+F_{1}+F_{2}+\sum_{r=3}^{p}F_{r}$
$\leq\binom{k}{2p}\binom{2p}{p}\|A\|_{2p}^{4p}+\binom{k}{2p-1}\binom{2p-1}{p}p\left(\|A\|_{2p}^{4p}+2\|A\|_{4p}^{4p}\right)$
$\quad+\binom{k}{2p-2}\binom{2p-2}{p}\binom{p}{2}\left(6\|A\|_{4p}^{4p}+3\|A\|_{4}^{4}\|A\|_{4p-4}^{4p-4}\right)$
$\quad+\sum_{r=3}^{p}\binom{k}{2p-r}\binom{2p-r}{p}\binom{p}{r}\sum_{\ell=1}^{r}c(r,\ell)\|A\|_{4}^{4(\ell-1)}\|A\|_{4p-4(\ell-1)}^{4p-4(\ell-1)}.\qed$
Proof of Corollary 5.
Note that $\mathrm{Var}(\hat{\theta}_{2p})=\binom{k}{p}^{-2}\mathrm{Var}(\hat{\Theta}_{2p})$. By Theorem 4 and the fact $\binom{k}{2p-r}\binom{2p-r}{p}\binom{p}{r}=\mathcal{O}(k^{2p-r})$, we have
$\mathrm{Var}(\hat{\theta}_{2p})$ $\leq-\|A\|_{2p}^{4p}+\binom{k}{p}^{-2}\binom{k}{2p}\binom{2p}{p}\|A\|_{2p}^{4p}$
$\quad+\binom{k}{p}^{-2}\binom{k}{2p-1}\binom{2p-1}{p}p\left(\|A\|_{2p}^{4p}+2\|A\|_{4p}^{4p}\right)$
$\quad+\binom{k}{p}^{-2}\binom{k}{2p-2}\binom{2p-2}{p}\binom{p}{2}\left(6\|A\|_{4p}^{4p}+3\|A\|_{4}^{4}\|A\|_{4p-4}^{4p-4}\right)+\mathcal{O}\left(\frac{1}{k^{3}}\right).$ (20)
For the second term in (20), performing a second-order expansion in $1/k$, we have
$\binom{k}{p}^{-2}\binom{k}{2p}\binom{2p}{p}\|A\|_{2p}^{4p}$ $=\frac{(k-p)(k-p-1)\cdots(k-2p+1)}{k(k-1)\cdots(k-p+1)}\|A\|_{2p}^{4p}$
$=\|A\|_{2p}^{4p}\prod_{i=0}^{p-1}\left(1-\frac{p}{k-i}\right)$
$=\left(1-\frac{p^{2}}{k}+\frac{p^{2}(p-1)(p-2)}{2k^{2}}\right)\|A\|_{2p}^{4p}+\mathcal{O}\left(\frac{1}{k^{3}}\right).$ (21)
For the third term in (20), using a first-order expansion of $\prod_{i=1}^{p-1}\left(1-\frac{p-1}{k-i}\right)$, we have
$\quad\binom{k}{p}^{-2}\binom{k}{2p-1}\binom{2p-1}{p}p\left(\|A\|_{2p}^{4p}+2\|A\|_{4p}^{4p}\right)$
$=\frac{p^{2}(k-p)(k-p-1)\cdots(k-2p+2)}{k(k-1)\cdots(k-p+1)}\left(\|A\|_{2p}^{4p}+2\|A\|_{4p}^{4p}\right)$
$=\frac{p^{2}}{k}\left(\|A\|_{2p}^{4p}+2\|A\|_{4p}^{4p}\right)\prod_{i=1}^{p-1}\left(1-\frac{p-1}{k-i}\right)$
$=\left(\frac{p^{2}}{k}-\frac{p^{2}(p-1)^{2}}{k^{2}}\right)\left(\|A\|_{2p}^{4p}+2\|A\|_{4p}^{4p}\right)+\mathcal{O}\left(\frac{1}{k^{3}}\right).$ (22)
For the fourth term in (20), we have
$\quad\binom{k}{p}^{-2}\binom{k}{2p-2}\binom{2p-2}{p}\binom{p}{2}\left(6\|A\|_{4p}^{4p}+3\|A\|_{4}^{4}\|A\|_{4p-4}^{4p-4}\right)$
$=\frac{p^{2}(p-1)^{2}}{2k(k-1)}\frac{(k-p)(k-p-1)\cdots(k-2p+3)}{(k-2)(k-3)\cdots(k-p+1)}\left(6\|A\|_{4p}^{4p}+3\|A\|_{4}^{4}\|A\|_{4p-4}^{4p-4}\right)$
$=\frac{p^{2}(p-1)^{2}}{2k^{2}}\left(6\|A\|_{4p}^{4p}+3\|A\|_{4}^{4}\|A\|_{4p-4}^{4p-4}\right)+\mathcal{O}\left(\frac{1}{k^{3}}\right).$ (23)
Finally, putting together (20)–(23), we have
$\mathrm{Var}(\hat{\theta}_{2p})$ $\leq-\|A\|_{2p}^{4p}+\left(1-\frac{p^{2}}{k}+\frac{p^{2}(p-1)(p-2)}{2k^{2}}\right)\|A\|_{2p}^{4p}$
$\quad+\left(\frac{p^{2}}{k}-\frac{p^{2}(p-1)^{2}}{k^{2}}\right)\left(\|A\|_{2p}^{4p}+2\|A\|_{4p}^{4p}\right)$
$\quad+\frac{p^{2}(p-1)^{2}}{2k^{2}}\left(6\|A\|_{4p}^{4p}+3\|A\|_{4}^{4}\|A\|_{4p-4}^{4p-4}\right)+\mathcal{O}\left(\frac{1}{k^{3}}\right)$
$\leq\frac{2p^{2}}{k}\|A\|_{4p}^{4p}+\frac{p^{2}(p-1)^{2}}{k^{2}}\left(\|A\|_{4p}^{4p}+\frac{3}{2}\|A\|_{4}^{4}\|A\|_{4p-4}^{4p-4}-\frac{1}{2}\|A\|_{2p}^{4p}\right).\qed$
4 Numerical examples
We numerically illustrate the performance of our bounds and estimates on the variance of $\hat{\theta}_{2p}$. The numerical experiments have been performed in Matlab R2022b, and all the figures in this document can be reproduced using the code available at https://github.com/Alice94/schatten-norm. Recall that, without loss of generality, we can work with diagonal matrices. In the plots below, we estimate the exact variance of $\hat{\theta}_{2p}$ for a fixed matrix and a fixed value of $k$ by running $\hat{\theta}_{2p}$ for $10000$ times and computing the sample variance of the results; this corresponds to the blue line in the figures in this section. The red lines (first-order estimate) correspond to Theorem 3, the yellow lines (second-order estimate) correspond to Theorem 5, and the purple dotted lines correspond to the bound of Theorem 4. In all our examples, we keep the matrix size quite small ($100\times 100$) so that we are able to run the estimator many times to compute its empirical variance as precisely as possible.
Example 15.
We consider the diagonal matrix $A\in\mathbb{R}^{100\times 100}$ with diagonal entries $0.8,0.8^{2},0.8^{3},\ldots,0.8^{100}$. Figure 3 illustrates the variance of $\hat{\theta}_{2p}$ for $p=4,6,8$ with values of $k$ ranging from $16$ to $500$. Note that the first-order estimate and the second-order estimates tend to underestimate the variance and get better for larger values of $k$, as expected.
Figure 3: Comparison of the variance of $\hat{\theta}_{2p}$, our bounds, and our estimates for the matrix from Example 15.
Example 16.
In Figure 4, we illustrate the variance of $\hat{\theta}_{2p}$ for $p=3,5,7$ for the $100\times 100$ diagonal matrix with diagonal entries $1,1/4,1/9,1/16,\ldots,1/10000$.
Figure 4: Comparison of the variance of $\hat{\theta}_{2p}$, our bounds, and our estimates for the matrix from Example 16.
Example 17.
The motivation for this work is to explore the behavior of $\hat{\theta}_{2p}$ for numerically rank matrices. However, even for the matrices with no eigenvalue decay, our bound from Theorem 4 improves the existing bound (2) from [6]. In Figure 5, we consider $A=I_{100}$, the identity matrix, and compare the empirical variance of $\hat{\theta}_{2p}$ with our estimates, our bounds, and the bound (2). Our bounds are not tight at all, especially for larger values of $p$, but are several orders of magnitude tighter than the bound (2).
Figure 5: Comparison of the variance of $\hat{\theta}_{2p}$, our bounds, our estimates, and bound (2) from [6] for the identity matrix.
5 Discussion
In this work, we analyzed the variance of $\hat{\theta}_{2p}$ given by Algorithm 1, which is an unbiased estimator for the $2p$-th power of the Schatten-$2p$ norm of a matrix $A$, assuming only a sketch $Y=A\Omega$ with a Gaussian matrix $\Omega$ is available. We provided a new bound – Theorem 4 – that improves the bound (2) from [6] for any matrix. Moreover, we focused on the matrices that have strong singular value decay, and proposed a first-order and second-order estimate on the variance. Numerically, we observe that these estimates are rather precise for moderate values of $p$ and $k$. For matrices with no singular value decay, the bound holds but the estimates do not really describe the behavior of $\hat{\theta}_{2p}$ since they are too optimistic.
The streaming setting we considered is applicable in scenarios where we need to extract information on the moments of a covariance matrix of a multivariate Gaussian distribution while only the samples drawn from the distribution is available. This is the original setting considered in [6]. Algorithm 1 was also suggested as a cheap Schatten-$2p$ norm estimator in the randomized numerical linear algebra review papers [8, 10]. We remark that, if the matrix $A$ can be accessed more than once via matrix-vector product, one should use other techniques to estimate Schatten norms; for instance, the ones based on Hutchinson trace estimator since they have smaller variance.